OpenStudy (anonymous):

I'll fan and medal to whoever assists me correctly first! ^_^ ~ Rewrite this equation in the form: y = ax + b y-3=1/2x+1 Thank you!! <3 xoxoxo

OpenStudy (anonymous):

me me pick me!!

OpenStudy (anonymous):

add 3 to both sides

OpenStudy (anonymous):

\[y-3=\frac{1}{2}x+1\] \[y=\frac{1}{2}x+1+3\]

OpenStudy (anonymous):

-3 cancels out, right?

OpenStudy (anonymous):

oh hehe

OpenStudy (anonymous):

yeah right make sure to add \(1+3\) before handing in your answer

OpenStudy (anonymous):

so y=1/2x+4 ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes of course